A ball of mass 625g traveling 4.00m/s


A ball of mass 625 g traveling 4.00 m/s to the right collides with another ball of equal mass that is initially at rest. The first ball is at rest after the collision. What is the velocity of the second ball after the collision?
Answer #1

Apply law of conservation of momentum

M1u1+M2u2= M1v1+M2v2

0.625×4+625×0= 0.625×0 + 0.625×v2

v2= 4m/s